OpenStudy (anonymous):

PLEASE I BEG OF YOU I WILL MEDAL AND FAN I BEG OF YOU PLEASE HELP Choose the missing number to make equivalent fractions. 7/BLANK=28/40 = A. 10 B. 19 C. 28 D. 160

OpenStudy (anonymous):

PLEASE HURRY

OpenStudy (anonymous):

DOH

OpenStudy (anonymous):

C I THANK

OpenStudy (chunkymonkay):

THE ANSWER IS A. 10

OpenStudy (anonymous):

it was a 10 but thanks anyway i will give you a medal

OpenStudy (anonymous):

lol im new how do i give medals

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

wait i have 2 more can you stay

OpenStudy (chunkymonkay):

click best resopnce

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (chunkymonkay):

thank you for the medal

OpenStudy (anonymous):

Which number can replace the variable to make these fractions equivalent? n/9 = 56/63 A. 2 B. 7 C. 8 D. 9

OpenStudy (anonymous):

times both 7/blank by 4 to get 28/40 and the answer is 10 so A

OpenStudy (anonymous):

and no prob

OpenStudy (anonymous):

i had a new guestion up

OpenStudy (chunkymonkay):

the answer is C

OpenStudy (anonymous):

C if you divide both by 7 the answer is 8/9 = 56/63